OpenStudy (anonymous):

Use the elimination method to solve the given system of equations. 5x +y=3 10x +2y=4 a.(0,3) b.(0,10) c.no solutions d.infinitely many solutions

OpenStudy (kropot72):

If (5x + y = 3) is multiplied by 2 on both sides it becomes 10x + 2y = 6. But the second equation says 10x + 2y = 4. Therefore there can be no solutions.

OpenStudy (anonymous):

Thanks Again :)

jhonyy9 (jhonyy9):

multiply first by -2 and assum those 5x+y=3 *(-2) -10x -2y =-6 10x +2y =4 -------------- 0 0 =-2 - so than result that the system have not solutions

OpenStudy (kropot72):

You're welcome:)